OpenStudy (fixxer):

I hope some experts out there could help me obtain the right plot for this solid. z=8-x-2y the plane is cut of by the cylinder with domain in the xy plane x^2+y^2=1

OpenStudy (fixxer):

I have tried to set \[8-x-2y=x^2+y^2\] this give me a parabolid and are faulty. Preferebly someone can tell me how to obtain a solution.

sam (.sam.):

Multivariable calculus? x,y,z plane?

OpenStudy (fixxer):

Yeah. \[r(t)=[r*\cos(t),r*\sin(t),plane]\] is that a possible solution? and what is the plane? 8-r*cos(t)-2*r*sin(t)? i wanna do it general in cartesian to

OpenStudy (fixxer):

is this correct?

OpenStudy (irishboy123):

yup, i think so intersection is on: \vec r = <x,y,z> = <x, y ,8 - x - 2y>\(\vec r = <x,y,z> = <x, y ,8 - x - 2y>\) = <\cot \theta,\sin \theta,8 - \cos \theta - 2\sin \theta>, \quad t \in [0, 2\pi ]\(= <\cot \theta,\sin \theta,8 - \cos \theta - 2\sin \theta>, \quad \theta \in [0, 2\pi ]\)